I have a problem setting up and solving word problems. Can someone please help me? Here is the word problem:
Cost of can A can in the shape of a right circular cylinder is required to have a volume of 500 cubic centimeters. The top and bottom are made of material that costs 6 cents per square centimeter, while the sides are made of material that costs 4 cents per square centimeter. Express the total cost C of the material as a function of the radius r of the cylinder. What will the cost be if the radius is 10 centimeters?

Answer by richwmiller(17219)

ends surface area=pi*r^2
ends area=pi*10^2= 100*pi=314
2*314
volume=pi*r^2*h
500=pi*10^2*h
h=5/pi
surface area
2pi*r*h=side area
2*pi*10*5/pi
side area =100
6*2*314= 3768 cost of top and bottom
4*100=400
3768+400=$41.68 total cost